1 pt) A P(X1126) Probability B. P(X < 966) Probability c. P(X > 1046) Probability
Answer #1

Solution:

n = 25

\mu = 1050

\sigma = 200

( A )

P( \bar{X} > 1126 ) =  \mu\pm z (\sigma / \sqrt{n} )

= P( z > ( 1126 - 1050 )/(200/\sqrt{25}) )

= P( z > 79 / 40 )

= P( z > 1.975 )

= 0.0239

( B )

P( \bar{X} < 966 )   =  \mu\pm z (\sigma / \sqrt{n} )

= P( z < ( 966 - 1050 )/(200/\sqrt{25}) )

= P( z < -2.1 )

= 0.0179

( C )

P( \bar{X} > 1046 ) =  =  \mu\pm z (\sigma / \sqrt{n} )

= P( z > ( 1046 - 1050 )/(200/\sqrt{25}) )

= P( z > -0.1)

= 0.5398